Use the Chain Rule to determine an equation for the acceleration when 𝑎 = 𝑑𝑣/ 𝑑𝑡

When 𝑣 = (2𝑡2 + 3)4

and

When 𝑣 = ln(4𝑡3 − 1)


1
Expert's answer
2022-01-04T18:07:40-0500

1)When "v = (2t^2+3)^4 = f(g(t))", where "f(t) = t^4, g(t) = 2t^2 + 3" .

Since "\frac{df}{dt} = 4t^3" and "\frac{dg}{dt} = 4t" , by the chain rule, we obtain that:


"a = \frac{dv}{dt} = \frac{df}{dg} \frac{dg}{dt} = 4(2t^2+3)^3 \cdot 4t = 16t(2t^2+3)^3"


2)When "v = ln(4t^3-1) = f(g(t))", where "f(t) =ln( 4t^3-1), g(t) = (4t^3 - 1)" .

Since "\frac{df}{dt} =\frac{1}{4t^3-1}" and "\frac{dg}{dt} =12t^2" , by the chain rule, we obtain that:


"a = \frac{dv}{dt} = \frac{df}{dg} \frac{dg}{dt} = \frac{1}{4t^3-1}\cdot 12t^2 = \frac{12t^2}{4t^3-1}"